Foster’s stocks one raw material item, Drum 4, which forms the raw material content for their sole product, and the information below relates to this stock item for May 2016.

Opening stock

       100 units                                                      Price $20 per unit

 

Receipts of stock item

2-5-19                                    250 units                    Price $25 per unit

13-5-19                                   300 units                    Price $30 per unit

21-5-19                                   220 units                    Price $35 per unit

 

Issues to production

 6-5-19                                 200 units

19-5-19                                   300 units

30-5-19                                   250 units

 

            Required:

(a)   For Drum 4, calculate the value of closing stock as at 31 May 2019, and the cost of materials issued to production for the month of May, using perpetual methods of inventory valuation for:

FIFO

LIFO and AVCO  




1
Expert's answer
2021-11-14T17:36:07-0500

Solution:

1.). FIFO method:

According to the FIFO method, the first units are sold first, and the calculation uses the newest units. 

Cost of material issued = (200 x 25) + (300 x 30) + (250 x 35) = 22,750

Closing stock = Opening stock + receipt of stock – Issues to production

Closing stock = 100 + 770 – 750 = 120 units

Value of closing stock = 120 x 20 = 2,400


2.). LIFO method:

According to LIFO method, determine the cost of your most recent inventory and multiply it by the amount of inventory sold.

Cost of material issued = (220 x 35) + (25 x 30) + (300 x 30) + (250 x 25) = 23,700


Closing stock = Opening stock + receipt of stock – Issues to production

Closing stock = 100 + 770 – 750 = 120 units

Value of closing stock = 120 x 35 = 4,200


3.). AVCO Method:

The weighted-average cost method takes the weighted average of all units in the company's inventory.

Total purchased = 770

Total issued = 750

Total closing inventory = 120


Weighted average cost:

250 x 25 = 6250

300 x 30 = 9000

220 x 35 = 7700

770        = 22,950

AVCO = 22,950/770 = 29.80

Cost of material issued = 750 x 29.80 = 22,350

Value of closing stock = 120 x 29.80 = 3,576